# Capturing strong correlation effects on a quantum annealer: calculation of avoided crossing in the H4 molecule using the quantum annealer eigensolver

### v1

@@ III Results and Discussions
IV Comparison of QAE and VQE
We devote this section to a qualitative comparison of the strengths and weaknesses of VQE, the leading NISQ algorithm for quantum chemistry, and the QAE algorithm. We add that we do not compare the performance of QAE with best known classical computing algorithms, since it is a NISQ approach, and a comparison with VQE is likely more suitable in that context. It is worth noting that we do not, for the purposes of the qualitative analysis, compare QAE with variants of VQE, but with only the traditional version of the latter.
We begin with general remarks on both the algorithms, before we discuss results from our numerical calculations of AC using the VQE algorithm.
  • The very first consideration is the absence of gates in QAE. Since gate errors are a major limiting factor in gate-based approaches of which VQE is one, the accuracy in results that one expects from a gate-based algorithm is usually limited, as opposed to approaches involving quantum annealing. The results from Ref. [5], where the authors achieve an accuracy of about 99 percent, indeed reflects this observation.
  • We also note that QAE possesses another inherent advantage in giving better results, since we pick the lowest energies across repetitions by construction. This is in contrast to VQE, where the mean expectation value across repetitions is typically chosen as the final result.
  • Furthermore, in VQE, the choice of ansatz is paramount. While physics-inspired/chemistry-based ansatze such as the unitary coupled cluster (UCC) based ones yield excellent results in noiseless simulations, the entire class of hardware efficient ansatze often suffer from the notorious barren plateau problem. On the other hand, we are unlikely to run into barren plateau issues in QAE, since by construction, we employ a CI wave function, which is physics-inspired/chemistry-based, and the algorithm involves an explore-exploit metaheuristic strategy, due to which we do not have a notion of iterations and convergence. It is worth adding that modified QAE workflows could have a convergence aspect to it, for example, see Ref. [5], but in our current study, we use the originally proposed QAE workflow, where we simply execute many shots of quantum annealing for each pre-selected $\lambda$ value within a range.
  • This also leads us to the next point: since VQE is iterative in nature, the convergence behaviour heavily relies on the choice of optimizer (with the correct choice not being necessarily easy), the precision sought, as well as the nature of the specific system of interest. On the other hand, one can think of quantum annealing itself as performing the optimization in the case of QAE.
  • The number of parameters to optimize grow identically for both UCC-based VQE and QAE (which relies on CI). For example, in a UCCSD (UCC in the singles and doubles approximation) ansatz-based VQE computation, the number of parameters to be optimized grow as $\sim n_{o}^{2}n_{v}^{2}$, where $n_{o}$ and $n_{v}$ refer to the number of occupied and unoccupied spin orbitals respectively. The number of parameters to optimize in QAE also grows identically for CISD (configuration interaction in the singles and doubles approximation).
  • The inputs to VQE are the one- and two-electron integrals. The inputs to QAE are the CI Hamiltonian (which in turn require one- and two- electron integrals) and also $\lambda$. The requirement of finding the latter adds an additional layer of complication to QAE.
  • We now move to the range of physical effects that the two approaches can capture. VQE-UCCSD, in its traditional form, is not suited to handle strong correlation effects of the type we address in this work, whereas QAE, which is based on CI, is. If VQE needs to address strong correlation effects, we need to employ suitable variants such as a multi-reference UCC (MRUCC) ansatz, which is typically accompanied by more quantum resources, or alternatives such as use of a double exponential UCC-like ansatz [41], etc. We also add here that comparing UCCSD-based VQE and FCI-based QAE is reasonable, especially since we are dealing with a small system in our work, and thus the difference in the quality of correlation effects captured by UCCSD and FCI are not as critical for the purposes of the analysis.
  • Both algorithms face some common challenges, such as the lack of a fully connected hardware can lead to incurring more quantum resources (more CNOT gates for VQE via SWAP gates, and more physical qubits for QAE via embedding). There exist some fully connected gate-based computers (for example, IonQ Aria and Forte devices, as well as the Quantinuum Model H1 and H2 machines), but we are yet to have such an annealing hardware (although work is underway on architectures such as the LHZ type [42]).
  • QAE also has the drawback of current QA hardware only being able to accommodate the transverse field Ising Hamiltonian, thus leading to encoding costs.
We now carry out the AC calculation in the VQE framework. We first calculate the ground state energy, followed by computing the excited state energy of interest to us in the quantum equation of motion VQE framework. We pick the AC geometry (2 Bohr). We use the Qiskit 1.3.1 package [43] throughout. We carry our 8-qubit, 26-parameter, noiseless VQE simulations with 1000 shots, and across 10 repetitions, we obtain a mean energy of -1.93288 Ha for the ground state energy. We recall that with 1000 shots and 10 repetitions, QAE, on the D-Wave hardware, yields a lower energy of -1.93381 Ha. As Fig. 8 shows, the mean energy obtained from a noiseless VQE simulation improves to -1.93875 Ha upon increasing the number of shots to 5000, which is better than the QAE estimate with 1000 shots and 10 repetitions. We note that the value is better than those obtained from QAE with 5000 shots and 10 repetitions (-1.93750 Ha) and with 5000 shots and 30 repetitions (-1.93760 Ha). We recall that the FCI value is -1.93943 Ha. Lastly, since VQE employs a single reference UCCSD ansatz, the results are obtained at the cost of slow convergence. The computation took 298 iterations to converge with the SLSQP optimizer, in contrast to 168 iterations that it took for convergence at 1.2 Bohr geometry and 113 at the 8 Bohr geometry, where strong correlation effects are less pronounced. A similar observation on a single reference ansatz yielding reasonable results at the cost of more iterations was observed in Ref. [13]. Our Quantum Equation of Motion (QEOM) state vector calculations [44] directly predicted an AC value of 0.14578 Ha (the excited state energy was -1.79250 Ha), which is within about 4 percent of the FCI AC value.
In order to place the QAE and VQE computations on an equal footing, we would also need to compute the ground state energy using VQE on real quantum hardware. We do not carry out such a calculation here, but assess the challenges that we anticipate in such a scenario. Assuming that we carry out the calculation on the IBM Brooklyn device, we incur 2244 $CX$ gates (the number of one-qubit gates are 3200, 1216, and 4 for $RZ$, $SX$, and $X$ respectively). We note that even if the device had been fully connected, the number of two-qubit gates is still 1440, which would yield a rather low result fidelity. For example, assuming an optimistic estimate of $0.999$ for the two-qubit gate fidelity, an estimate of the result fidelity even with a fictitious all-to-all connected topology is a mere $\sim 0.999^{1440}\approx 0.2$. Thus, in order to obtain meaningful results, one needs to aggressively optimize the quantum circuit in addition to implementing additional resource reduction strategies to reduce the number of measurements, etc [22]. The authors show in Ref. [22] that with a suite of such classically resource-intensive strategies, the number of gates can be reduced to the extent that makes computation on current-day quantum devices possible.
@@ IV Comparison of QAE and VQE
Figure 8: Comparison of the ground state energy ($E_{g}$) obtained using VQE across different number of shots. For each shot number, we repeat the VQE calculation 10 times (10 repetitions). The HF and the FCI values are provided for reference.

### v4

@@ III.1 QAE results for avoided crossings
III.2 VQE results for ground state energies
For our VQE calculations, we only consider the ground state energy, although it is possible in principle to obtain the excited state energy of interest, using an equation of motion based VQE approach. QAE naturally incorporates CI allowing it to accurately predict energies both near equilibrium and at far-from-equilibrium geometries, effectively capturing both weak and strong correlation effects. In contrast, VQE with the UCCSD ansatz works best for near-equilibrium geometries and is unable to account for strong correlation effects. A fair comparison with QAE would require that we employ a multi-reference theory, for example, the icMRUCC (internally contracted multi-reference unitary coupled cluster) [46], where the reference state is a linear combination of multiple determinants. However, pursuing a multi-reference treatment poses certain challenges. First, one must survey the class of multi-reference theories and identify the one that is best suited for VQE in terms of cost of implementation. Furthermore, in multi-reference theories where redundancies are not naturally accounted for, the computational cost to remove redundancies (duplicate states that need to be removed) is not necessarily easy. For example, in the icMRUCC theory, the process of obtaining the indices for the linearly independent excitation amplitudes involves computing square root of an overlap matrix, which is computationally hard for large matrices.
Given the current limitations in two-qubit gate fidelities, we implement only a single-parameter VQE using the UCCSD ansatz, focusing on measuring only the dominant terms in the Hamiltonian. To capture correlation energy effectively in a multi-reference approach, we would need to measure at least two dominant terms from the Hamiltonian. Consequently, the multi-reference theory requires more resources. Thus, our UCCSD results serve as an upper bound to the values obtained from a multi-reference UCCSD approach, which, if implemented on a quantum computer with improved gate fidelities, would yield better quality results for the ground state energy than UCCSD in strongly correlated regimes.
We carry out our 26-parameter VQE-UCCSD calculation on the ibm_kiev device. The choice of hardware was made by comparing the CLOPS for the three currently available quantum computers (Brisbane, Sherbrooke, and Kyiv), and picking the one with the highest CLOPS. All of our calculations were performed with 4096 shots, and we quote our result by averaging over 10 repetitions. We begin by specifying the system parameters: processor type: Eagle R3, native gates: {ECR, ID, RZ, SX, X}, median ECR error: 1.086×$\times 10^{-2}$, median readout error: 1.367×$\times 10^{-2}$, median T1: 289.88 $\mu$s, median T2: 113.13 $\mu$s, and connectivity: heavy hexagonal lattice topology. We incur 1716 $ECR$ gates (the number of one-qubit gates are 16500, 5844, and 1444 for $RZ$, $SX$, and $X$ respectively). We note that even if the device had been fully connected, $\mathcal{N}_{2qg}$ is still 1440, which would yield a rather low result fidelity. For example, assuming an optimistic estimate of $0.999$ for the two-qubit gate fidelity (for comparison, state-of-the-art commercial quantum computers such as the IonQ Forte-I have current fidelities hovering in and around $0.99$), an estimate of the result fidelity even with a fictitious all-to-all connected topology is a mere $\sim 0.999^{1440}\approx 0.2$. Thus, in order to obtain meaningful results, one needs to aggressively optimize the quantum circuit in addition to implementing additional resource reduction strategies to reduce the number of measurements, etc [15]. The authors of the cited work show that with a suite of such classically resource-intensive strategies, the number of gates can be reduced to the extent that makes computation on current-day quantum devices possible. We borrow their techniques for our computations, and briefly describe them below before delving into our findings from the computations themselves. We begin by carrying out a VQE calculation on a traditional computer. We then carry out optimization on two fronts: wave function (the quantum circuit itself) and Hamiltonian (reducing the number of terms in the Hamiltonian, so that the number of circuits evaluated is reduced). On the circuit optimization side, we first pass our quantum circuit with the optimal parameters into three optimization routines sequentially: Qiskit $(L=3)$ [47], Pytket [48], and PyZX [49]. This was followed by an additional layer of circuit optimization using reinforcement learning-based ZX calculus routine and a causal flow preserving ZX calculus module [50], followed by another round of Qiskit $(L=3)$ optimization to yield a quantum circuit with the following gate count: $RZ$: 38, $SX$: 19, $X$=11, and $ECR$: 12. On the Hamiltonian front, we group the Hamiltonian into cliques, which are sets of mutually (qubit-wise) commuting terms, and then group cliques that yield common energies into supercliques, based on ideas from Ref. [51]. We then pick the dominant superclique (in energy) for our hardware computation.
We measured two circuits (that belong to the two dominant supercliques) on the optimized quantum circuit built with optimized parameters to obtain the ground state energy. We find that in spite of reducing $\mathcal{N}_{2qg}$ to 12, we obtain an energy of $-1.71197$ Ha (recalling that they were obtained by averaging over 10 repeats, each with 4096 shots), with a PFD of 7.81 percent relative to the value obtained in the same setting on a simulator (noiseless). It is also worth adding that the correlation energy itself in the noiseless setting is $-0.08015$ Ha, whereas the one predicted from the IBM execution is significantly different at 0.06480 Ha. Augmenting the computations with error mitigation techniques (outside the scope of the current work) can improve the quality of the result, but it is clear that in order to carry out more resource intensive MRUCC-VQE computations even in small active spaces such as the one we have chosen would require further advances on the hardware front. On the other hand, the QAE algorithm gives reasonable values for the ground state energy ($-1.93760$ Ha with 5000 shots and averaged over 30 repetitions) on the D-Wave hardware.
@@ III.2 VQE results for ground state energies
Figure 8: Comparison of the ground state energy ($E_{g}$) obtained using VQE across different number of shots. For each shot number, we repeat the VQE calculation 10 times (10 repetitions). The HF, the FCI, and the QAE (5000 shots and averaged over 30 repetitions) values are provided for reference.
IV Comparison between scaling behaviours of QAE and VQE
In this section, we carry out a preliminary comparison of QAE and VQE, and wherever relevant, comment on scaling. Our analysis is preliminary in the following sense: We consider only the chemistry-inspired unitary coupled cluster ansatz and limited to the singles and doubles approximation (UCCSD) for VQE, and determinant-based configuration interaction in the singles and doubles approximation (CISD) for QAE (we use GUGA-based CI in our numerical calculations in the previous sections, but consider determinant-based CI for the scaling analysis on grounds of simplicity; as and when possible, we comment on GUGA-based CI). Furthermore, we do not compare QAE with variants of VQE, but with only the traditional version  [1] of the latter. We also do not attempt to address in depth the open (yet critical) questions from both algorithms such as input state preparation in the case of VQE and anneal time scaling in the case of QAE. Figure 9 presents schematics of both the algorithms, where common numerical labels indicate points of comparison.
Refer to caption Refer to caption
(a) (b)
Figure 9: Schematic of the QAE (left panel) and the VQE (right panel) algorithms.
IV.0.1 Problem inputs
  • One- and two- electron integrals: Both QAE and VQE require us to supply one- and two- electron integrals. They are obtained from a classical computer, involving the HF (scales typically as $N^{4}$ for $N$ spin-orbitals [52]) and atomic orbital to molecular orbital integral transformation (scales as $N^{5}$ [52]) steps. The number of Hamiltonian terms themselves scale as $N^{4}$. This overhead can be reduced by employing techniques such as Hamiltonian factorization [53, 54, 55]. Reducing this overhead has the effect of reducing the number of circuit evaluations in VQE per iteration (since there are as many circuits per iteration as the number of Hamiltonian terms). On the other hand, in QAE, the ‘factorization’ needs to be performed on the wave function in order to reduce the number of the quantum annealing runs. In fact, the GUGA-based approach that we employ in this work corresponds to this procedure. The determinant-based wave function expansion is inefficient with respect to the number of variables, because Slater determinants are not necessarily spin eigenfunctions. Qualitatively, GUGA generates spin symmetry-adapted CSFs by block-diagonalizing the electron spin $\mathbf{S}^{2}$ operator to reduce the number of variables to optimize.
  • Evaluating Hamiltonian matrix elements: Following the computation of integrals, the CI Hamiltonian matrix elements must be constructed classically prior to performing QAE. We therefore analyze the cost associated with such a construction. If there are $d_{CI}$ coefficients, then noting that the Hamiltonian is hermitian, we need to evaluate $\frac{d_{CI}^{2}}{2}+\frac{d_{CI}}{2}\sim d_{CI}^{2}$ elements. We consider the CISD approximation for the purpose of the analysis. We also introduce some standard notation: $n_{o}$ refers to the total number of occupied spin orbitals (i.e. total number of electrons), and $n_{v}$ the number of unoccupied spin orbitals. A cursory look shows that since the number of matrix elements to evaluate goes as $n_{o}^{4}n_{v}^{4}$ and the cost of constructing a CI matrix element, given the one- and two- electron integrals, by using the Slater–Condon rules goes at most as $n_{o}^{2}$ (corresponding to the case of the wavefunction in the bra and ket being the same), the cost of obtaining the CI Hamiltonian matrix elements is at most $n_{o}^{6}n_{v}^{4}$, that is, $N^{10}$. However, this naive evaluation does not account for sparsity of the Hamiltonian (via the Slater–Condon rules) and also assumes that every matrix element evaluation is as costly as the costliest one’s evaluation. A more careful approach by accounting for the aforementioned considerations indicates a much cheaper scaling of $N^{6}$. The details of the derivation are presented in Appendix B. The scaling behavior for evaluating matrix elements in GUGA-CI is very similar to that of determinant-based approach. However, GUGA-CI incurs an extra cost which comes from evaluating additional coefficients associated with pathways of the Shavitt graph, which can introduce an extra scaling of $\mathcal{O}(N)$. Therefore the overall computational cost for GUGA scales as $\mathcal{O}(N^{7})$ [56, 57]. While the number of Hamiltonian matrix elements are fewer in GUGA-based CI than for determinant-based CI, the time taken to evaluate them can be longer for the former, due to evaluation of the aforementioned coefficients. Having discussed the cost of constructing the matrix elements, we now move onto analyzing the encoding cost of the energy functional in equation (2) to that given by equation (4).
    Refer to caption Refer to caption
    (a) (b)
    Figure 10: Optimisation domain for $N=2$ and $K=2$ in (a) the space of binary variables and (b) the space of real variables.
  • Encoding in QAE in QUBO form: We address the question of the scaling of the number of CI coefficients, $c_{i}$, as well as required scaling for the parameter, K.
    • Although the number of coefficients in FCI scale as $d_{CI}\sim\binom{n_{o}}{m}\binom{n_{v}}{m}\sim\frac{n_{o}^{m}n^{m}_{v}}{(m!)^{2}}$, where $m$ is the excitation level, one performs truncated CI calculations where $m$ is fixed (for example, CISD, where $m=2$), and thus the number of coefficients, in practice, scales polynomially as $N^{4}$. We recall that K represents the number of qubits used to discretise each coefficient such that the optimisation problem is over a set of binary variables. We qualitatively examine the scaling of K with the number of coefficients, $d_{CI}$. Each coefficient $c_{i}$ is approximated by a K bit binary representation defined as
      $c_{i}\sim\tilde{c}_{i}=0.c_{i1}\ldots c_{ij}\ldots c_{iK},$
      $\forall j\text{ and each }i=1,2,\ldots,d_{CI}$
      where $c_{ij}\in\{0,1\}$ implying that each coefficient can take $2^{K}$ distinct values. This allows us to visualise the binary optimisation domain as an $d_{CI}$ dimensional hyper-cube with each side consisting of $2^{K}$ discrete points. For simplicity we show the case for $d_{CI}=2$ and $K=2$ in Figure 10 (a). The resulting space is uniformly discretised with neighbouring points spaced $\frac{1}{2^{K}}$ distance apart. Furthermore, we consider a unit cell of this space, which is a hyper-cube of volume $\frac{1}{2^{Kd_{CI}}}$ and circumscribe it by a hyper-sphere, as seen in Figure 10(b). Thus, the discretization error is upper bounded by the radius of this hyper-sphere. Assuming equal discretization $\Delta x$ in each dimension, we have
      $r^{2}$ $=\sum^{d_{CI}}_{i=1}\bigg{(}\frac{\Delta x}{2}\bigg{)}^{2}$ (5)
      $r$ $=\sqrt{d_{CI}}\frac{\Delta x}{2}$
      For constant error we demand $\sqrt{d_{CI}}\sim\frac{1}{\Delta x}$. Since $\Delta x=\frac{1}{2^{K}}$ we have
      $K\sim\log_{2}d_{CI}$ (6)
      Therefore K must grow logarithmically with the number of coefficients, that is, as $log_{2}{d_{CI}}$ which is equal to $log_{2}N^{4}$ for CISD. We now consider the corresponding cost in VQE, namely that for input state preparation via a quantum cicuit.
    Figure 11: Figure illustrating the growth of the number of two qubit gates, $\mathcal{N}_{2qg}$, with number of qubits (written as $N$ in the figure), for input state preparation in VQE.
  • Input state preparation in VQE: On the other hand, in VQE, one needs to construct the input reference state via an isometry, which in the case of a multi-reference UCC method, can be non-trivial. We note that not only does one need to employ classical many-body methods to find the coefficients of the entangled state, but also the cost of an input state preparation isometry in general scales exponentially in the number of two-qubit gates, $\mathcal{N}_{2qg}$ [58]. Using chemical intuition to guess a small number of determinants (for instance, assuming a log log growth of the number of determinants in the number of all possible determinants) reduces the scaling ($\lceil log(log(2^{N}))\rceil\sim\lceil log(N)\rceil$), but the resulting circuits are still very deep for the NISQ era. We consider a toy example where we go from $N$ of 4 to 16 in steps of 2, and thus the number of determinants for the input state would go from 2 to 4. We restrict the number of occupied spin-orbitals to 2, and find that $\mathcal{N}_{2qg}$ increases as $2^{N}$ (the choice for coefficients is explained in Appendix C, while Figure 11 presents the result from our numerical analysis as a plot), in spite of a $loglog$ reduction in the number of determinants (the worst case scaling for $2^{N}$ basis states is $4^{N}$ two-qubit gates [59, 58]). This unsurprising conclusion even for a limited toy example underscores the hardness of one of the biggest open problems in gate-based algorithms: efficient input state preparation. Having said this, we also stress that our conclusions are drawn from numerical simulations that employ Qiskit’s isometry routines, and there are works in literature that attempt to address the topic of manual construction of polynomial-depth quantum circuits to prepare multi-configurational wave functions (for example, see Refs. [60] and  [61]).
  • The Lagrange mutiplier: In QAE, we need to select a range of $\lambda$ values to scan, as well as the number of values in a selected range. This choice is not obvious, especially in light of the results shown in Figure 7, and we defer this analysis to a future study.
IV.0.2 The main processing steps
For VQE, this corresponds to the cost function evaluation on a quantum computer and parameter updates on a classical computer. For QAE, this corresponds to the quantum annealing step.
  • Choice of ansatz: In VQE, the choice of ansatz is paramount to convergence. While physics-inspired/chemistry-based ansatze such as the UCC based ones yield excellent results in noiseless simulations, the entire class of hardware efficient ansatze often suffer from the notorious barren plateau problem [62]. On the other hand, we are unlikely to run into barren plateau issues in QAE, since by construction, we employ a CI wave function, which is physics-inspired/chemistry-based, and the algorithm involves an explore-exploit metaheuristic strategy, due to which we do not have a notion of iterations and convergence. It is worth adding that modified QAE workflows could have a convergence aspect to it, for example, see Ref. [17], but in our current study, we use the originally proposed QAE workflow, where we simply execute many shots of quantum annealing for each pre-selected $\lambda$ value within a range.
  • Number of two-qubit gates in ansatz: A very important consideration is the absence of gates in QAE. Since gate errors are a major limiting factor in gate-based approaches of which VQE is one, the accuracy in results that one expects from a gate-based algorithm is usually limited, as opposed to approaches involving quantum annealing. The results from Ref. [17], where the authors achieve an accuracy of about 99 percent, indeed reflects this observation. On the other hand, in VQE with the UCCSD ansatz for example, the number of two-qubit gates scale as $\sim N^{5}$. We can arrive at the estimate by considering that a Pauli gadget over $N$ qubits (recalling that the number of spin-orbitals, $N$, is the number of qubits in VQE) scales at most linearly in $N$ (assuming the Jordan–Wigner transformation), and that there are $n_{o}^{2}n_{v}^{2}$ double excitations and thus as many Pauli words. The analysis assumes order one step one Trotterization. Thus, for example, even for a calculation involving a small molecule, where $n_{o}$ and $n_{v}$ are about 20 and 40 respectively, $\mathcal{N}_{2qg}$ is already of the order of 100 million, and even with aggressive quantum circuit optimization strategies (for example, see Ref. [15]), it is unlikely to carry out such calculations at the physical qubit level on current and near-future quantum computers.
  • Initial guess for VQE parameters: In VQE, we need to choose the initial guess for the parameters. Although we choose zeroes as initial guess values for our parameters, one can accelerate convergence by feeding in better-informed guesses such as MP2 (Møller–Plesset perturbation theory to second order in energy) initial guess (for example, see Ref. [6]). This, however, is accompanied by appropriate classical pre-processing steps.
  • Parameter update step in VQE: In VQE, the parameter update limits prospects of achieving speed-up in large system sizes. For our numerical simulations whose results we will discuss in Section III.2, we use the SLSQP optimizer, which is widely employed in literature. The cost function evaluation per iteration scales quadratically in the number of parameters (Refer section 2.2.4 in [63]). This implies the cost scales roughly as $\sim N^{8}$. However, since the scaling of the number of iterations with system size is highly problem-specific and non-trivial in nature, we defer the analysis to a future study.
  • Number of parameters to optimize: This grows identically for both UCC-based VQE and QAE (which relies on CI). For example, in a UCCSD ansatz-based VQE computation, the number of parameters to be optimized grow as $n_{o}^{2}n_{v}^{2}\sim N^{4}$. However, it is important to also note that in the VQE-UCC framework, some higher-order excitations are accounted through the non-linear expansion of the UCC state, and the number of Slater determinants (or CSFs as in our case) included in the wave function expansion is larger for VQE-UCC than for QAE, even if the same number of parameters are included. For instance, even though UCCSD explicitly considers only $T_{1}$ and $T_{2}$ terms, it inherently includes correlation contributions from terms like $T_{3},T_{1}^{2},T_{1}^{3},T_{1}^{4},T_{1}T_{2}$ etc. Therefore, while the number of parameters remain unchanged, the number of determinants involved is significantly larger in UCCSD compared to CISD.
  • Convergence: This also leads us to the next point: since VQE is iterative in nature, the convergence behaviour heavily relies on the choice of optimizer (with the correct choice not being necessarily easy), the precision sought, as well as the nature of the specific system of interest. On the other hand, one can think of quantum annealing itself as performing the optimization in the case of QAE.
  • We note that QAE possesses another inherent advantage in giving better results, since we pick the lowest energies across repetitions by construction. This is in contrast to VQE, where the mean expectation value across repetitions is typically chosen as the final result. This is because evaluating the expectation value of an observable $A$ using VQE involves computing the weighted sum $\sum_{i}\lambda_{i}p_{i}$, where $\{\lambda_{i}\}$ are the eigenvalues of $A$. Here $p_{i}=\tilde{N}_{i}/N_{s}$ represents the probability of the state being in the $i^{th}$ eigenstate, $\tilde{N}_{i}$ denotes the number of shots yielding that eigenstate, and $N_{s}$ is the total number of shots. Assuming the exact distribution has probabilities given by $p^{0}_{i}=\tilde{N_{i}^{0}}/N_{s}$, achieving an accurate estimate of the expectation value requires generating sufficient statistics such that $\tilde{N}_{i}\approx\tilde{N}^{0}_{i}$, $\forall i$. In contrast, in QAE, a single occurrence of the minimum energy is sufficient in order for it to qualify as the optimal solution.
  • Number of shots: The previous point leads us to the question of the number of shots required for a calculation. If we seek a precision $\epsilon$ in our calculation, VQE requires that we supply $\sim N^{4}/\epsilon^{2}$ number of shots (see Ref. [64], Section 5.1.1). The cost function evaluation in each iteration and for each Hamiltonian term involves cost from number of shots and the number of two-qubit gates for that circuit, which as we saw scales as $\sim N^{5}$. On the other hand, for QAE, determining the scaling behaviour of the number of shots is highly non-trivial, although it is central to the completion of the scaling analysis. Furthermore, noting that the anneal time itself scales as the inverse of minimum gap squared (where gap refers to the energy gap between the ground state and the first excited state of the total Hamiltonian, whose final Hamiltonian is the QUBO described in Eq. 4), we require that for annealing to be efficient, the gap must scales polynomially in $N$. We defer these two important considerations of the anneal time and the number of shots scaling behaviours with system size for a future work, but note that they may favourably or adversely impact the overall scaling of QAE.
IV.0.3 Embedding
  • We now discuss embedding. The number of coefficients to optimize post QAE encoding is $n_{coeff}=d_{CI}\times K$ which results in an $n_{coeff}$ qubit fully connected problem graph. Minor embedding an $n_{coeff}$ qubit all-to-all connected problem requires a physical qubit overhead which in the worst case scales as $\mathcal{O}(n^{2}_{coeff})$ [65]. Therefore, the number of physical qubits scale as
[…]
    In the gate model, the overhead of embedding can be approximately quantified in terms of the additional number of SWAP gates required to account for the limited connectivity between physical qubits. Consider the worst-case scenario of an $N$ qubit quantum circuit where every pair of qubits is connected by a CNOT. Suppose the hardware has only nearest neighbour 1D connectivity, that is, the qubits are arranged as nodes on a path graph (worst case scenario). In this case, the total number of CNOT gates which cannot be executed directly due to connectivity constraints are $\bigg{(}\frac{N(N-1)}{2}-N\bigg{)}$. Therefore the total number of SWAP gates incurred, $\mathcal{N}_{\mathrm{additional\ SWAPs}}$ are bounded by
    $\mathcal{N}_{\mathrm{additional\ SWAPs}}$ (8)
    $\leq\bigg{(}\frac{N(N-1)}{2}-N\bigg{)}\bigg{(}N-2\bigg{)}$
    which in the worst case scales as $\mathcal{O}(N^{3})$. As a consequence, the overall result fidelity ($F_{\mathrm{res}}$) would degrade by a factor of $\sim F^{\mathcal{N}_{\mathrm{additional\ SWAPs}}}$ where $F$ denotes the two qubit gate fidelity. This scaling arises from
    $F_{\mathrm{res}}\sim F^{\mathcal{N}_{\mathrm{2qg}}+\mathcal{N}_{\mathrm{additional\ SWAPs}}}.$ (9)
    resulting in an exponential decrease of the result fidelity.
  • In VQE, although there are gate-based quantum computers to execute the algorithm with over 100 qubits, we are limited by the number of gate operations even for small numbers of physical qubits. On the other hand, in QAE, we are limited by the number of physical qubits. For example, we find that for a given repetition, that is, across hundred $\lambda$ values that are scanned for our chosen $\lambda$ range, the number of physical qubits can vary between about 700 to about 900 to construct our QUBO with 80 logical qubits. This also reflects when we wish to increase K in a computation. For example, were we to carry out our computations on H4 with a split valence basis set such as the 6-31G (4 electron- 5 orbital active space), a QUBO size of 170 (assuming $K=10$) requires on an average about 4000 physical qubits (over those scanned $\lambda$ values for which an embedding was found). Finding an embedding with any further increase in K is not possible, since the D-Wave annealers have only 5000 physical qubits. In such cases, one needs to opt for Sub-QUBOs (for example, see Ref. [66]), and additional errors incurred in approximations involved in choice of sub-QUBOs also add to the error budget. It is worth adding that the embedding issue we consider is specific to current state-of-the-art D-Wave hardware and not to the QAE algorithm itself; future advances in this direction, for example, machines with substantially better connectivity, could alleviate embedding-related issues.
Table 2: Table summarizing the worst case scaling for various steps of the QAE (for determinant-based CISD) and the VQE (for a multi-reference UCC) algorithms. The quantum computing overheads are marked in bold. In the table, $N$ is the total number of spin-orbitals.
Steps QAE (end-to-end) VQE (end-to-end)
Hartree-Fock $N^{4}$ $N^{4}$
One- and two- electron integrals $N^{5}$ $N^{5}$
Evaluation of $H_{ij}$ $N^{6}$-
Number of CI coefficients $N^{4}$-
K scaling $\log_{2}(N)$-
$\bm{\mathcal{N}^{\textbf{iso}}_{\text{{2qg}}}}$- $2^{N}$ [^1] Under the assumption that the number of determinants grow as $\sim log(N)$, for an $N$ qubit state for a representative example. Furthermore, our estimates are based on Qiskit’s isometry routine.
$\bm{\mathcal{N}^{\textbf{uccsd}}_{\textbf{2qg}}}$- $N^{5}$
Optimizer -- $N^{8}$ [^2] Cost per iteration for the SLSQP optimizer.
Embedding Num physical qubits (CISD) $\leq N^{8}log_{2}^{2}(N)$ Num additional SWAPs $\leq N^{3}$
IV.0.4 Net scaling behaviours
Table 2 summarizes the upper bounds for the individual steps of the QAE and the VQE algorithms (end-to-end). Since the entire analysis is preliminary, we take a simplistic route and assume that the costs can be summed, and thus we pick the highest scaling term for the scaling of an algorithm. We immediately see that the input state preparation step in VQE is exponentially costly (this is not the case for single reference UCC-VQE, since the input state preparation only requires one-qubit gates), and only under the assumption that an oracle supplies the input state does the optimizer step becomes the most steeply scaling. On the other hand, in QAE, embedding is the costliest step. Thus, QAE scales as $N^{8}log_{2}^{2}(N)$ and VQE as $N^{8}$ per iteration, with the latter assuming the SLSQP optimizer. For comparison, a regular execution of a CCSD or a CISD routine on a traditional computer scales as $N^{6}$. In view of our various simplifications (for instance, VQE: SLSQP is not the optimizer of choice outside statevector type calculations, and QAE: we assume a $log(log(2^{N})$ behaviour for the growth of number of determinants in $N$), our finding does not definitely establish the scaling for QAE and VQE, and is to only be viewed as a first step towards a more rigorous scaling analysis for the future. Furthermore, VQE and QAE are both restricted by the current state-of-the-art quantum hardware; even with the current best NISQ computers, it is impractical to go beyond 5000 physical qubits and about 50 two-qubit gates for QAE and VQE executions respectively, and thus we do not expect any advantage anyway over traditional computers for such limited system sizes.